# Graph Alignment via Dual-Pass Spectral Encoding and Latent Space Communication

### v1

@@ C.4.3 Vision and language models
We adopt the set of 32 vision models used in Blind Match [Schnaus et al., 2025]. For self-supervised methods, we use DINO [Caron et al., 2021] models (RN50 and ViT-S/B with patch sizes 16 and 8) trained on ImageNet-1k and DINOv2 [Oquab et al., 2023] models (ViT-S/B/L/G with patch size 14) trained on the LVD-142M dataset, as well as fully supervised models such as DeiT variants [Touvron et al., 2021] (Tiny, Small, and Base with patch size 16, including distilled and high-resolution @384 versions) and ConvNeXt models[Liu et al., 2022b] (Base and Large, pretrained on ImageNet-1k or ImageNet-22k, with additional fine-tuned @384 variants). For vision–language pretraining, we employ CLIP [Ramesh et al., 2022] with both ResNet backbones (RN50, RN101, RN50x4, RN50x16, RN50x64) and Vision Transformer architectures (ViT-B/32, ViT-B/16, ViT-L/14, ViT-L/14@336). All experiments are conducted using official implementations and pretrained weights, ensuring consistent and reliable representation extraction for each model.
@@
We consider four pretrained language models spanning diverse architectures and training paradigms, including the RN50x4 model from CLIP [Ramesh et al., 2022] and three models—all-MiniLM-L6-v2, all-mpnet-base-v2, and all-Roberta-large-v1—extracted from the SentenceTransformers library [Reimers and Gurevych, 2019]. All vision and language models used in our experiments are summarized in Table 6.
@@
DINO [Caron et al., 2021]: RN50, ViT-S/16, ViT-S/8, ViT-B/16, ViT-B/8
DINOv2 [Oquab et al., 2023]: ViT-S/14, ViT-B/14, ViT-L/14, ViT-G/14
DeiT [Touvron et al., 2021]: DeiT-T/16, DeiT-T/16d, DeiT-S/16, DeiT-S/16d, DeiT-B/16, DeiT-B/16@384, DeiT-B/16d, DeiT-B/16d@384
ConvNeXt [Liu et al., 2022b]: CN-B-1, CN-B-22, CN-L-1, CN-L-22, CN-L-22ft@384, CN-XL-22ft@384
CLIP [Ramesh et al., 2022]: RN50, RN101, RN50x4, RN50x16, RN50x64, ViT-B/32, ViT-B/16, ViT-L/14, ViT-L/14@336
@@
CLIP [Ramesh et al., 2022]: RN50x4
SentenceTransformers [Reimers and Gurevych, 2019]: all-MiniLM-L6-v2, all-mpnet-base-v2, all-Roberta-large-v1
Appendix D Ablation study on graph encoders
We evaluate the impact of different GNN encoder architectures on the alignment accuracy within the GADL framework using two real-world benchmark datasets. Specifically, we conduct a comparative evaluation of our proposed dual-pass GCN encoder against four GNN variants: GCN [Kipf and Welling, 2017], GIN [Xu et al., 2018a], JKGNN [Xu et al., 2018b], and TIDE [Behmanesh et al., 2023].
For these encoders, we adopt a 6-layer architecture with ReLU activation functions, following the configurations presented in their respective papers. Additionally, TIDE is applied in a single-channel setup, where the learnable parameter $t$ is shared across all channels.
@@ Appendix D Ablation study on graph encoders
Appendix E Hyperparameter sensitivity
We conduct a sensitivity analysis to examine the influence of key hyperparameters on model performance. In particular, we focus on the loss weighting coefficients $\lambda_{\text{FM}}$, $\lambda_{\text{bij}}$, and $\lambda_{\text{orth}}$, which control the relative importance of the functional map loss, bijectivity constraint, and orthogonality regularization, respectively, in the overall training objective.
@@ Appendix E Hyperparameter sensitivity
Appendix F Comprehensive results for vision-language model combinations
To provide a comprehensive evaluation of vision–language alignment, we test our proposed model across multiple vision and language model combinations on two benchmark datasets, CIFAR-10 and CINIC-10. The vision models considered include CLIP [Ramesh et al., 2022], ConvNeXt [Liu et al., 2022b], DINO [Caron et al., 2021], DINOv2 [Oquab et al., 2023], and DeiT [Touvron et al., 2021]. For the language models, we include the RN50x4 model from CLIP [Ramesh et al., 2022] as well as three models from the SentenceTransformers library [Reimers and Gurevych, 2019]: all-MiniLM-L6-v2, all-mpnet-base-v2, and all-Roberta-large-v1.
@@ Appendix F Comprehensive results for vision-language model combinations
These results indicate that our proposed approach consistently achieves higher matching accuracies across diverse vision–language encoder combinations, outperforming state-of-the-art baselines such as the Hahn-Grant solver in most configurations (cf. Figure 4 in the Hahn-Grant paper [Schnaus et al., 2025]).
@@
We further evaluate the proposed model across diverse vision–language combinations on the larger-scale CIFAR-100 and ImageNet-100 benchmarks. Table 7 summarizes the results.These experimental results demonstrate that model performance is highly context-dependent, with no single architecture achieving universal superiority across CIFAR-100 and ImageNet-100.

### v2

@@ C.5.2 Vision and language models
We adopt the set of 32 vision models used in Blind Match (Schnaus et al., 2025). For self-supervised methods, we use DINO (Caron et al., 2021) models (RN50 and ViT-S/B with patch sizes 16 and 8) trained on ImageNet-1k and DINOv2 (Oquab et al., 2023) models (ViT-S/B/L/G with patch size 14) trained on the LVD-142M dataset, as well as fully supervised models such as DeiT variants (Touvron et al., 2021) (Tiny, Small, and Base with patch size 16, including distilled and high-resolution @384 versions) and ConvNeXt models(Liu et al., 2022b) (Base and Large, pretrained on ImageNet-1k or ImageNet-22k, with additional fine-tuned @384 variants). For vision–language pretraining, we employ CLIP (Ramesh et al., 2022) with both ResNet backbones (RN50, RN101, RN50x4, RN50x16, RN50x64) and Vision Transformer architectures (ViT-B/32, ViT-B/16, ViT-L/14, ViT-L/14@336). All experiments are conducted using official implementations and pretrained weights, ensuring consistent and reliable representation extraction for each model.
@@
We consider four pretrained language models spanning diverse architectures and training paradigms, including the RN50x4 model from CLIP (Ramesh et al., 2022) and three models, all-MiniLM-L6-v2, all-mpnet-base-v2, and all-Roberta-large-v1, extracted from the SentenceTransformers library (Reimers & Gurevych, 2019). All vision and language models used in our experiments are summarized in Table 5.
@@
DINO (Caron et al., 2021): RN50, ViT-S/16, ViT-S/8, ViT-B/16, ViT-B/8
DINOv2 (Oquab et al., 2023): ViT-S/14, ViT-B/14, ViT-L/14, ViT-G/14
DeiT (Touvron et al., 2021): DeiT-T/16, DeiT-T/16d, DeiT-S/16, DeiT-S/16d, DeiT-B/16, DeiT-B/16@384, DeiT-B/16d, DeiT-B/16d@384
ConvNeXt (Liu et al., 2022b): CN-B-1, CN-B-22, CN-L-1, CN-L-22, CN-L-22ft@384, CN-XL-22ft@384
CLIP (Ramesh et al., 2022): RN50, RN101, RN50x4, RN50x16, RN50x64, ViT-B/32, ViT-B/16, ViT-L/14, ViT-L/14@336
@@
CLIP (Ramesh et al., 2022): RN50x4
SentenceTransformers (Reimers & Gurevych, 2019): all-MiniLM-L6-v2, all-mpnet-base-v2, all-Roberta-large-v1
Appendix D Ablation analysis
To analyze the impact of individual components in the proposed framework, we conduct an ablation study evaluating variants with specific modules removed or modified. We compare GADL against: (1) GADL w/o dual-pass encoder: replaces the dual-pass GCN with a standard single-pass GCN while retaining latent-space communication; (2) GADL w/o bijectivity regularization: removes the bijectivity regularization term while keeping the dual-pass encoder; (3) GADL w/o orthogonality regularization: omits orthogonality regularization while maintaining all other components.
Results are summarized in Table 6. They highlight the individual contribution of each component to the overall alignment performance. Replacing the dual-pass GCN with a standard encoder causes substantial accuracy drops. Furthermore, the removal of the regularization terms notably degrades performance, confirming their importance in achieving robust and reliable node alignment.
Method ACM-DBLP Douban Online-Offline
Hit@1 Hit@5 Hit@10 Hit@50 Hit@1 Hit@5 Hit@10 Hit@50
GADL 88.63 94.76 96.16 98.41 53.31 73.61 80.67 94.18
Appendix E Ablation study on graph encoders
We evaluate the impact of different GNN encoder architectures on the alignment accuracy within the GADL framework using two real-world benchmark datasets. Specifically, we conduct a comparative evaluation of our proposed dual-pass GCN encoder against four GNN variants: GCN (Kipf & Welling, 2017), GIN (Xu et al., 2018a), JKGNN (Xu et al., 2018b), and TIDE (Behmanesh et al., 2023). For these encoders, we adopt a 6-layer architecture with ReLU activation functions, following the configurations presented in their respective papers. Additionally, TIDE is applied in a single-channel setup, where the learnable parameter $t$ is shared across all channels.
@@ Appendix E Ablation study on graph encoders
Appendix F Hyperparameter sensitivity
We conduct a sensitivity analysis to examine the influence of key hyperparameters on model performance. In particular, we focus on the loss weighting coefficients $\lambda_{\mathrm{FM}}$, $\lambda_{\mathrm{bij}}$, and $\lambda_{\mathrm{orth}}$, which control the relative importance of the functional map loss, bijectivity constraint, and orthogonality regularization, respectively, in the overall training objective.
@@ Appendix F Hyperparameter sensitivity
Appendix G Comprehensive results for vision-language model combinations
To provide a comprehensive evaluation of vision–language alignment, we test our proposed model across multiple vision and language model combinations on two benchmark datasets, CIFAR-10 and CINIC-10. The vision models considered include CLIP (Ramesh et al., 2022), ConvNeXt (Liu et al., 2022b), DINO (Caron et al., 2021), DINOv2 (Oquab et al., 2023), and DeiT (Touvron et al., 2021). For the language models, we include the RN50x4 model from CLIP (Ramesh et al., 2022) as well as three models from the SentenceTransformers library (Reimers & Gurevych, 2019): all-MiniLM-L6-v2, all-mpnet-base-v2, and all-Roberta-large-v1.
@@ Appendix G Comprehensive results for vision-language model combinations
These results indicate that our proposed approach consistently achieves higher matching accuracies across diverse vision–language encoder combinations, outperforming state-of-the-art baselines such as the Hahn-Grant solver in most configurations (cf. Figure 4 in the Hahn-Grant paper (Schnaus et al., 2025)).
@@
We further evaluate the proposed model across diverse vision–language combinations on the larger-scale CIFAR-100 and ImageNet-100 benchmarks. Table 7 summarizes the results. These results demonstrate that model performance is highly context-dependent, with no single architecture achieving universal superiority across CIFAR-100 and ImageNet-100.